# Friederike Becker1, Fabian Krüger1 and Melanie Schienle1,2

### v1

@@ 3.2 Choice of Tuning Parameters
Tables 4 and Figure 6 in the appendix summarize the training sample performance of various combinations of the tuning parameters. Based on these results, and the considerations in Section 2, we use a rolling window of $R=11$ observations, the absolute error method, and the default empirical quantile method (type = 7) [^1] Note that for $R=11$ and the absolute error method, this is also equivalent to the inverse of the empirical distribution function, that is, type = 1. As we do not estimate quantiles at the tails of the distribution the arguments of Hyndman and Fan (1996) in favor of the type = 8 method do not apply for our setting. of the quantile function in R. Our training sample results in Figure 6 indicate that shorter window lengths $R<11$ perform worse in terms of the IS. The difference in performance is particularly pronounced for very short window lengths $74\leq R<$ 7, and at longer forecast horizons. By contrast, the benefit of increasing $R$ from $8$ to $11$ is remarkably small in all empirical setups (absolute and directional errors, all forecast horizons). Hence the marginal benefit of increasing the sample size seems to vanish rapidly as $R$ increases. We exclude choices $R>11$ since we cannot evaluate their performance given the time span of our training data, while an extension of the training data would have reduced the time span of our holdout sample, thus limiting our ability to assess out-of-sample performance. From a statistical perspective, using a short window length $R$ ensures adaptability to potential structural breaks (c.f. Inoue et al., 2017). We find that coverage rates on the training data are generally more favorable for the absolute than for the directional error method, sometimes on the order of about 10 percentage points (see Table 4), while interval scores are similar.
@@
Figure 2: Illustration of the process that we use to construct prediction intervals at the $50\%$ and $80\%$, using forecasts and realized values from the International Monetary Fund World Economic Outlook forecast data.
@@ 3.3 Benchmark Models
In order to put the performance of our proposed method in perspective, we compare it to three benchmarks. For the first benchmark (denoted AR), we compute point predictions from an autoregressive model of order one, estimated via ordinary least squares. We fit the model to quarterly data on the predictand (inflation or GDP growth) obtained from the OECD, and construct a forecast of the annual series as a weighted sum of quarterly forecasts. See Section B.1 for details. We then construct prediction intervals from empirical quantiles of the model’s past forecast errors, analogous to our proposed procedure for the IMF forecasts.
For the second benchmark (denoted BVAR), we consider a Bayesian vector autoregressive model specification proposed by Primiceri (2005) and Del Negro and Primiceri (2015), and implemented in the R package bvarsv (Krüger, 2015). In our main results, we consider two quarterly system variables (inflation and GDP growth) obtained from the OECD. In a robustness check described below, we add an index of financial stress as a third variable. Section B.1 provides details. As for the IMF forecasts (and for the AR model), we then construct prediction intervals from empirical quantiles of the BVAR’s past forecast errors.
@@
The AR and BVAR benchmarks are analogous to our proposed methodology (based on IMF forecasts), except that they use a different point forecast. Hence they allow to assess the competitiveness of the IMF’s point forecasts. While AR is simple and transparent, BVAR is much more flexible, allowing for time-varying parameters. The latter comes at the cost of increased complexity (in particular, choice of prior parameters and simulation of the posterior distribution based estimation via Markov chain Monte Carlo methods). In contrast to AR and BVAR, BVAR-direct is fully based on parametric modeling assumptions.
@@ 3.4 Forecast Performance on the Hold-Out Data
For inflation, IMF-based forecasts tend to perform best at the shortest horizon (fall, current year), whereas scores are otherwise similar for all models. The decomposition into the score components is also similar, with dispersion accounting for a sizeable portion of the penalties. Empirical coverage values tend to stay close to nominal levels, for all methods.
@@
Figure 5 shows the average length of the prediction intervals of IMF-extracted probabilistic forecasts, separately for each combination of target variable and forecast horizon. For both inflation and real GDP growth, lengths shrink considerably between next-year and same-year horizons, as well as between the two same-year horizons. The reduction in size is less considerable between the two next-year horizons, especially for inflation, suggesting that the average level of confidence stays mostly similar from one-and-a-half years to one year out from the target. This result is in accordance with Qu et al. (2019), who find no significant evidence that the WEO’s average forecast accuracy improves during this time period.
@@ 3.6 Outlook: Prospective Real-Time Forecasts and Evaluation
Initial releases of the forecasts for October 2023 and April 2024 are available in a public GitHub repository (Becker et al., 2023), ensuring transparency and accountability through time stamps. The latter repository also links to the website https://probability-forecasting.shinyapps.io/macropi/ on which we provide concise visualizations of all forecasts and prediction intervals. Moreover, we plan to deposit a preregistration protocol (Becker et al., 2023) in which we make a commitment regarding how we will evaluate forecasts in 2029, after collecting five years of prospective forecasts and prediction intervals.
@@ Appendix A Uncertainty Quantification in Central Bank Reports
We are searching for concretely quantified statements about the uncertainty surrounding macro-economic forecasts in the G7’s central banks’ “flagship” reports for communicating their growth and inflation projections. For this, we applied the following protocol.
We note that this protocol does not necessarily provide a comprehensive view about central banks’ handling of uncertainty quantification and that each central bank might issue explicitly quantified statements about the uncertainty surrounding their macroeconomic forecasts elsewhere. The protocol thus only provides a snapshot view of whether and how central banks communicate uncertainty in their “flagship” reports for macroeconomic projections.
@@
      Given a match, we determined from context whether the given statement contained any uncertainty quantification for the macroeconomic forecasts in the document.
@@ B.2 Training Sample Performance
Table 4: Interval scores and interval coverage rates in the training sample (2001-2012), for the directional and the absolute error extraction method. In each block, that is, separately for each combination of target series, horizon, and evaluation metric, the better performing method is highlighted, .
Figure 6: Interval scores and empirical coverage rates in the training sample (2001-2012), for both targets and error extraction methods, and for various choices of the rolling window length $R$. Interval scores are averaged for the $50\%$ and $80\%$ interval. The column header indicates the target and error extraction method.
@@ B.3 Choice of Truth Source
Figure 8: Same as Figure 3, but using OECD truth values as the outcome data.
@@
Figure 9: Same as Figure 4, but using OECD truth values as the outcome data.
@@ B.4 Including a Financial Stability Index into the BVAR Benchmark
Figure 10: Same setup as Figure 3, but with the two BVAR specifications. ‘BVAR-direct’ is the same method as in Figure 3, whereas ‘BVAR-CISS’ additionally includes data from the CISS financial stability index. Here we exclude Japan and Canada from scoring, as we lack predictions from the BVAR - CISS model for these countries.

### v2

@@ 3.2 Choice of Tuning Parameters
Table 4 and Figure 6 in the appendix summarize the training sample performance of various combinations of the tuning parameters. Based on these results, and the considerations in Section 2, we use a rolling window of $R=11$ observations, the absolute error method, and the default empirical quantile method (type = 7) [^1] Note that for $R=11$ and the absolute error method, this is also equivalent to the inverse of the empirical distribution function, that is, type = 1. As we do not estimate quantiles at the tails of the distribution the arguments of Hyndman and Fan (1996) in favor of the type = 8 method do not apply for our setting. of the quantile function in R. Our training sample results indicate that shorter window lengths $R<11$ perform worse in terms of the IS, see Figure 6 in the appendix. The difference in performance is particularly pronounced for very short window lengths $74\leq R<$ 7 and at longer forecast horizons. By contrast, the benefit of increasing $R$ from $8$ to $11$ is remarkably small in all empirical setups (absolute and directional errors, all forecast horizons). Hence the marginal benefit of increasing the sample size seems to vanish rapidly as $R$ increases. We exclude choices $R>11$ since we cannot evaluate their performance given the time span of our training data, while an extension of the training data would have reduced the time span of our holdout sample, thus limiting our ability to assess out-of-sample performance. From a statistical perspective, using a short window length $R$ ensures adaptability to potential structural breaks (c.f. Inoue et al., 2017). We find that coverage rates on the training data are generally more favorable for the absolute than for the directional error method, sometimes on the order of about 10 percentage points (see Table 4 in the appendix), while interval scores are similar.
@@
Figure 2: Illustration of the process that we use to construct prediction intervals at the $50\%$ and $80\%$ confidence levels, using point forecasts and realized values from the International Monetary Fund World Economic Outlook forecast data.
@@ 3.3 Benchmark Models
In order to put the performance of our proposed method in perspective, we compare it to three benchmarks. For the first benchmark (denoted AR), we compute point predictions from an autoregressive model of order one, estimated via ordinary least squares. We fit the model to quarterly data on the predictand (inflation or GDP growth) obtained from the OECD, and construct a forecast of the annual series as a weighted sum of quarterly forecasts. See Section B.1 in the appendix for details. We then construct prediction intervals from empirical quantiles of the model’s past forecast errors, analogous to our proposed procedure for the IMF forecasts.
For the second benchmark (denoted BVAR), we consider a Bayesian vector autoregressive model specification proposed by Primiceri (2005) and Del Negro and Primiceri (2015), and implemented in the R package bvarsv (Krüger, 2015). In our main results, we consider two quarterly system variables (inflation and GDP growth) obtained from the OECD. In a robustness check described below, we add an index of financial stress as a third variable. Section B.1 in the appendix provides details. As for the IMF forecasts (and for the AR model), we then construct prediction intervals from empirical quantiles of the BVAR’s past forecast errors.
@@
The AR and BVAR benchmarks are analogous to our proposed methodology (based on IMF forecasts), except that they use a different point forecast. Hence they allow to assess the competitiveness of the IMF’s point forecasts in our setup. While AR is simple and transparent, BVAR is much more flexible, allowing for time-varying parameters. The latter comes at the cost of increased complexity (in particular, choice of prior parameters and simulation of the posterior distribution based estimation via Markov chain Monte Carlo methods). In contrast to AR and BVAR, BVAR-direct is fully based on parametric modeling assumptions.
@@ 3.4 Forecast Performance on the Hold-Out Data
For inflation, IMF-based forecasts tend to perform best at the shortest horizon (fall, current year), whereas scores are otherwise similar for all models. The decomposition into the score components is also similar, with dispersion and underprediction accounting for a sizeable portion of the penalties. Empirical coverage values tend to stay close to nominal levels, for all methods.
@@
Figure 5 shows the average length of the prediction intervals of IMF-extracted probabilistic forecasts, separately for each combination of target variable and forecast horizon. For both inflation and real GDP growth, lengths shrink considerably between next-year and same-year horizons, as well as between the two same-year horizons. The reduction in size is less considerable between the two next-year horizons, especially for inflation, suggesting that the average level of confidence stays mostly similar from one-and-a-half years to one year out from the target. This result is in accordance with Qu et al. (2019), who don’t find consistent significant evidence that the WEO’s average forecast accuracy improves during this time period.
@@ 3.6 Outlook: Prospective Real-Time Forecasts and Evaluation
Initial releases of the forecasts for October 2023 and April 2024 are available in a public GitHub repository (Becker et al., 2023), ensuring transparency and accountability through time stamps. The latter repository also links to the website https://probability-forecasting.shinyapps.io/macropi/ on which we provide concise visualizations of all forecasts and prediction intervals. Moreover, we have deposited a preregistration protocol (Becker et al., 2024) in which we make a commitment regarding how we will evaluate forecasts in 2029, after collecting prospective forecasts and prediction intervals for five years.
@@ Appendix A Uncertainty Quantification in Central Bank Reports
We are searching for concretely quantified statements about the uncertainty surrounding macroeconomic forecasts in the G7’s central banks’ “flagship” reports for communicating their growth and inflation projections. For this, we applied the following protocol.
We note that this protocol does not necessarily provide a comprehensive view about central banks’ handling of prediction uncertainty and that each central bank might issue explicitly quantified statements about the uncertainty surrounding their macroeconomic forecasts elsewhere. The protocol thus only provides a snapshot view of whether and how central banks communicate uncertainty in their “flagship” reports for macroeconomic projections.
@@
      Given a match, we determined from context whether the given statement contained any quantification of prediction uncertainty for the macroeconomic forecasts in the document.
@@ B.2 Training Sample Performance
This section includes results for the competing methods in the training sample (2001-2012).
@@
Table 4: Interval scores and interval coverage rates in the training sample (2001-2012), for the directional and the absolute error extraction method. In each block, that is, separately for each combination of target series, horizon, and evaluation metric, the better performing method is highlighted.
Figure 6: Interval scores and empirical coverage rates in the training sample (2001-2012), for both targets and error extraction methods, and for various choices of the rolling window length $R$. Interval scores are a weighted average of the $50\%$ and $80\%$ interval. The column header indicates the target and error extraction method.
@@ B.3 Choice of Truth Source
Figure 8: Same as Figure 3 in main paper, but using OECD truth values as the outcome data.
@@
Figure 9: Same as Figure 4 in main paper, but using OECD truth values as the outcome data.
@@ B.4 Including a Financial Stability Index into the BVAR Benchmark
Figure 10: Same setup as Figure 3 in main paper, but with the two BVAR specifications. ‘BVAR-direct’ is the same method as in Figure 3 in main paper, whereas ‘BVAR-CISS’ additionally includes data from the CISS financial stability index. Here we exclude Japan and Canada from scoring, as we lack predictions from the BVAR - CISS model for these countries.